Using the definition of the scalar product, find t
Answer #1

A.B = mod(A)*mod(B)*cos(thetha)

a) 6*8+7*5 = sqrt(6^2+7^2)*sqrt(8^2+5^2)*cos(theta)

=> cos(theha) = 0.954

thetha= 17.44 degree

b)

-3*3-3*4 = sqrt(3^2+3^2)*sqrt(3^2+4^2+2^2)*cos(theta)

=> cos(theha) = -0.9191

thetha = 156.79 degree

c)

-2*3+2*4 = sqrt(1^2+2^2+2^2)*sqrt(3^2+4^2)*cos(theta)

=> cos(theha) = 0.1333

thetha = 82.33 degree